OpenStudy (anonymous):

A forest ranger spots a fire from a 28-foot tower. The angle of depression from the tower to the fire is 11. To the nearest foot, how far is the fire from the base of the tower? Show the steps you use to find the solution.

OpenStudy (anonymous):



OpenStudy (anonymous):

1. \[\tan(\theta)=\frac{ x }{ 28 }\] 2.\[\tan(11º)=\frac{ x }{ 28 }\] 3. \[28\tan(11º)=x\]

OpenStudy (anonymous):

btw 28tan(11) is approximately: 5.443 feet

OpenStudy (anonymous):

So the final answer would be: 5 FEET

OpenStudy (anonymous):

tan is opposite over adjacent

OpenStudy (anonymous):

Wait...so how would i put this into a short answer?

OpenStudy (anonymous):

@mihirb

OpenStudy (anonymous):

Ok so basically say that the tan of the angle of depression is the distance from the fire over the height of the tower. and then say you solved for the distance from the fire

OpenStudy (anonymous):

Okay thanks :)